OpenStudy (anonymous):

Simplify. (4^3)^5 4^-2 4^8 4^15

OpenStudy (welshfella):

use the rule of indices (x^a)^b = x^ab

OpenStudy (welshfella):

the exponents are multiplied

OpenStudy (anonymous):

i dont get it..

OpenStudy (anonymous):

\[(a^b)^c=a^{bc}\]

OpenStudy (anonymous):

oh okay so it would be the last option then? 4^15?

OpenStudy (anonymous):

using this property simplify. \[(4^{3})^5=4^{3*5}=4^15\]

OpenStudy (welshfella):